# Joint identification of spatially variable genes via a network-assisted Bayesian regularization approach

### v1

@@ Abstract
Identifying genes that display spatial patterns is critical to investigating expression interactions within a spatial context and further dissecting biological understanding of complex mechanistic functionality. Despite the increase in statistical methods designed to identify spatially variable genes, they are mostly based on marginal analysis and share the limitation that the dependence (network) structures among genes are not well accommodated, where a biological process usually involves changes in multiple genes that interact in a complex network. Moreover, the latent cellular composition within spots may introduce confounding variations, negatively affecting identification accuracy. In this study, we develop a novel Bayesian regularization approach for spatial transcriptomic data, with the confounding variations induced by varying cellular distributions effectively corrected. Significantly advancing from the existing studies, a thresholded graph Laplacian regularization is proposed to simultaneously identify spatially variable genes and accommodate the network structure among genes. The proposed method is based on a zero-inflated negative binomial distribution, effectively accommodating the count nature, zero inflation, and overdispersion of spatial transcriptomic data. Extensive simulations and the application to real data demonstrate the competitive performance of the proposed method.
Keywords: Spatial transcriptomic data; Network analysis; Bayesian analysis
@@ 1 Introduction
For complex diseases, the recently developed and rapidly advancing spatial transcriptomics (ST) technology has provided a powerful tool to identify cellular organizations of gene expressions in tissues, which have long been recognized to be intimately linked to critical biological functions (Rao et al., 2021). One crucial task in ST data analysis is to identify genes with varying expressions across space, termed as spatially variable (SV) genes (Svensson et al., 2018). SV genes have been demonstrated to be associated with disease characteristics such as immune cell infiltration and tumor cell proliferation (Zuo et al., 2024), hence facilitating the discovery of tumorigenesis mechanisms and the development of therapeutic strategies.
@@
Many approaches have been recently proposed for SV gene detection. A great majority of them are based on the Gaussian process (GP). For example, SpatialDE (Svensson et al., 2018) models normalized expression data via GP regression and tests the significance of the spatial covariance matrix for each gene separately. SPARK (Sun et al., 2020), BOOST-GP (Li et al., 2021), and GPcounts (BinTayyash et al., 2021) also take advantage of GP regression but directly model raw count data with the Poisson, Negative Binomial (NB), and zero-inflated NB (ZINB) distributions, respectively. CTSV (Yu and Luo, 2022), on the other hand, implements a slightly different technique based on ZINB regression where the mean expression level is parameterized as a linear combination of functions of spatial coordinates. There are some non-parametric approaches with more computational efficiency, such as SPARK-X (Zhu et al., 2021) based on covariance-based testing, MERINGUE (Miller et al., 2021) based on Voronoi tessellation and classical Moran’s I score, and HEARTSVG (Yuan et al., 2023) based on constant variance testing. Despite considerable successes, the results of the aforementioned works are still sometimes unsatisfactory due to the high dimensionality of genes, high levels of noise and sparsity, and low resolution of spots. In addition to the transcriptomic data, other biological information, such as cellular phenotypes and genetic interactions, is often available and potentially provides a valuable complement to the present analysis. The integration of such assisted information into the transcriptomic analysis is a promising direction to mitigate the aforementioned challenges in ST data and further advance existing ST research.
@@
Specifically, first, most existing approaches for SV gene detection rely on marginal analysis, which models each gene separately and has the limitation that the dependence among genes is not well utilized. Increasing evidence has shown that diseases are mostly a result of a combination of multiple genetic alterations, and genetic factors usually interact with each other and are involved in a biological network (Barrio-Hernandez et al., 2023) (Figure 1). Specifically, genes connected within a network are believed to have similar biological functions, leading to potentially similar contributions to cellular organizations and functional mechanisms. Recently, an increasing number of biological networks have been amassed, such as protein-protein interaction (PPI) networks, metabolic networks, and regulatory networks. Curated network information has been widely adopted as a powerful supplement to gene expression analysis, particularly in bulk and single-cell sequencing analysis (Elyanow et al., 2020; Qin et al., 2023). However, the integration of network information for SV gene detection remains limited.
@@
Second, despite prosperous developments in recent years, measurements obtained using the sequencing-based ST technologies, such as Slide-seq and 10X Genomics Visium, are still “spot”-based, where the gene expression measurement at a single spot is usually a mixture of diverse cells from heterogeneous types rather than at a single-cell resolution (Yu and Luo, 2022) (Figure 1). This cellular composition diversity among spots has been demonstrated to probably contribute to expression variability (Cable et al., 2022), potentially confounding the detection of biologically relevant SV genes. In the published studies, there are very few approaches that conduct simultaneous cellular composition accommodation and SV gene detection. Limited existing studies include SPARK-X, which converts the inferred major cell types into binary indicators to mitigate the confounding effect caused by latent cell type distributions. Alternatively, CTSV concentrates on the detection of cell-type-specific SV genes whose expressions are affected by the spatial coordinates of cells of the same type.
@@
Motivated by the aforementioned challenges, we propose a Bayesian regularization approach for the joint identification of SV genes as shown in Figure 1. Specifically, the ZINB distribution is adopted to account for the count nature, sparsity, and overdispersion of raw expression measurements. The mean-value-based strategy is applied for SV gene detection, which enjoys the advantages of simplicity, interpretability, and computational efficiency. The most substantial advancement is that the proposed approach conducts joint detection with the network dependency structure among genes well incorporated, making it a big step forward from the existing marginal analysis. Moreover, the proposed approach introduces a series of cell-type-specific parameters to effectively correct for the confounding variations induced by varying cell-type compositions across spots. Overall, this study provides a practically useful and biologically meaningful approach for SV gene identification in ST analysis, with improved performance over alternatives as demonstrated in both simulation studies and application to a real-world ST dataset.
Figure 1: Overview.
@@ 2.1 Network and Cellular Composition Information for Integration
Consider an undirected network $G(V,E)$ that is constructed using biological information, where $V$ is the node set consisting of $p$ genes and $,E=\left\{e(j,l),j,l\in\{1,\right.$$\left.\cdots,p\}\right\}$ is the set of edges between nodes. For genes connected within the network, it is expected that they have similar biological functionalities, leading to potentially similar spatial variability, and thus are more likely to be SV or non-SV genes simultaneously. To induce such network-assisted identification, an adjacency matrix $\boldsymbol{A}=(a_{jl})_{p\times p}$ is first constructed based on $G(V,E)$, with $a_{jl}=1$ if there is an edge $e(j,l)$ between the $j$th and $l$th genes and $a_{jl}=0$ otherwise, and ${a}_{jj}=0$, for $j,l=1,2,\ldots,p$.
As for cellular compositions, denote $\boldsymbol{w}_{i}=\left(w_{i1},\dots,w_{iK}\right)^{\mathrm{T}}$ as the vector of cellular proportions for the $i$th spot, which satisfies the constraint that $0\leq w_{ik}\leq 1,k=1,\dots,K$ and $\sum_{k=1}^{K}w_{ik}=1$. Such information is typically available as ground truth or can be obtained using deconvolution methods such as RCTD (Cable et al., 2022), Redeconve (Zhou et al., 2023), and SONAR (Liu et al., 2023). Since the distributions of cellular proportions usually display spatial relatedness, which may confound SV gene detection, we correct for this potential confounding in SV gene identification.
@@ 2.2 Network-assisted Bayesian Modeling
We adopt the ZINB model due to its superiority for simultaneously accommodating the count measure, over-dispersion, and excessive zeros caused by dropouts. For SV gene identification, significantly different from the previous studies that introduce spatial differential expression via a covariance matrix, which is usually space- and time-consuming and lacks efficiency and scalability for large-scale analysis, we adopt the mean-value-based formulation. The spatial variability is intuitively reflected by $\beta^{(1)}_{j}$ and $\beta^{(2)}_{j}$, where the genes with $\beta^{(1)}_{j}\neq 0$ or $\beta^{(2)}_{j}\neq 0$ are regarded as SV. This requires less storage space and also involves a simpler estimation procedure. Moreover, we innovatively utilize a set of cell-type-specific $\alpha_{k}$’s for eliminating the confounding impact of the latent cellular composition, where spots that are spatially closer are often observed to have similar cell-type proportions (Cable et al., 2022). Different from the mean-value-based strategy adopted by Yu and Luo (2022) for marginal cell-type-specific SV gene detection, we focus on the detection of global SV genes while accommodating the cell-type proportion confounding.
@@ 2.3 Priors Specification
[…]
@@
The proposed priors have been motivated by the following considerations. The identification of SV genes is achieved using the hard-thresholded Gaussian prior, where $\beta^{(d)}_{j}$ is shrunk to zero when $\mathrm{t}_{\lambda,\rho^{(d)}_{j}}\left(\gamma^{(d)}_{j}\right)=\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\cdot\rho^{(d)}_{j}\right)=0$ ($d=1,2$). The thresholded Gaussian prior has been shown as a useful alternative to shrinkage priors in Bayesian sparse analysis (Wu et al., 2024), and is favored here for its simplicity and flexibility as well as its appealing interpretability as a minimal detectable signal. Here, we further introduce a series of weights $\rho_{j}^{(d)}$’s in $\mathbf{t}_{\lambda,\boldsymbol{\rho}}(\boldsymbol{\gamma})$ to adjust the shrinkage of various $\gamma^{(d)}_{j}$’s to improve selection efficiency, where the genes with strong spatial variability are potentially assigned small weights and thus are more likely to have nonzero $\beta^{(d)}_{j}$’s.
@@
Moreover, the network dependency is introduced via the graph Laplacian matrix in the covariance matrix of the hard-thresholded Gaussian prior. The proposed network-assisted strategy is motivated by the successes of Laplacian shrinkage in high-dimensional regression analysis (Chakraborty and Lozano, 2019; Cai et al., 2020). Different from these studies, we innovatively conduct SV gene detection with the network structures among genes incorporated. In particular, the Laplacian matrix $\widetilde{\mathbf{L}}$ is further modified with a pre-defined sign matrix $\operatorname{sgn}(\tilde{\boldsymbol{\beta}})\operatorname{sgn}(\tilde{\boldsymbol{\beta}})^{\mathrm{T}}$ to accommodate the scenario where two neighborhood genes are negatively correlated and have opposite directions of spatial variability. With the proposed priors, we have $\boldsymbol{\gamma}^{\mathrm{T}}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)\boldsymbol{\gamma}=\sum_{j\sim l}\left(\frac{\text{sgn}\left(\tilde{\beta}_{j}^{(1)}\right)\gamma_{j}^{(1)}}{\sqrt{d_{j}}}-\frac{\text{sgn}\left(\tilde{\beta}_{l}^{(1)}\right)\gamma_{l}^{(1)}}{\sqrt{d_{l}}}\right)^{2}+\sum_{j\sim l}\left(\frac{\text{sgn}\left(\tilde{\beta}_{j}^{(2)}\right)\gamma_{j}^{(2)}}{\sqrt{d_{j}}}-\frac{\text{sgn}\left(\tilde{\beta}_{l}^{(2)}\right)\gamma_{l}^{(2)}}{\sqrt{d_{l}}}\right)^{2}$, where for genes $j$ and $l$ with an edge $e(j,l)$, the absolute values of $\gamma^{(d)}_{j}$ and $\gamma^{(d)}_{l}$ are promoted to be similar, further inducing simultaneous SV or non-SV.

### v2

@@ Abstract
Identifying genes that display spatial patterns is critical to investigating expression interactions within a spatial context and further dissecting biological understanding of complex mechanistic functionality. Despite the increase in statistical methods designed to identify spatially variable genes, they are mostly based on marginal analysis and share the limitation that the dependence (network) structures among genes are not well accommodated, where a biological process usually involves changes in multiple genes that interact in a complex network. In addition, the latent cellular composition within the spots can introduce confounding variations, negatively affecting the accuracy of the identification. In this study, we develop a novel Bayesian regularization approach for spatial transcriptomic data, with confounding variations induced by varying cellular distributions effectively corrected. Significantly advancing from existing studies, a thresholded graph Laplacian regularization is proposed to simultaneously identify spatially variable genes and accommodate the network structure among genes. The proposed method is based on a zero-inflated negative binomial distribution, effectively accommodating the count nature, zero inflation, and overdispersion of spatial transcriptomic data. Extensive simulations and applications to real data demonstrate the competitive performance of the proposed method.
Bayesian regularization,
keywords:
[…]
, , and ,
t1corresponding author: wu.mengyun@mail.shufe.edu.cn t2the co-first authors
@@ 1 Introduction
Recently developed and rapidly advancing spatial transcriptomics (ST) technology enables gene expression profiling across numerous spatial locations within a tissue, offering biological insights into various contexts (Rao et al., 2021), with particularly prominent applications in cancer research. One crucial task in ST data analysis is to identify genes with varying expressions across space, termed as spatially variable (SV) genes (Svensson, Teichmann and Stegle, 2018). SV genes have been shown to be associated with disease characteristics such as immune cell infiltration and tumor cell proliferation (Zuo, Xia and Chen, 2024), thus facilitating the discovery of tumorigenesis mechanisms and the development of therapeutic strategies.
@@
Many approaches have recently been proposed for the detection of SV genes. The great majority of them are based on the Gaussian process (GP). For example, SpatialDE (Svensson, Teichmann and Stegle, 2018) models normalized expression data using GP regression and tests the significance of the spatial covariance matrix for each gene separately. SPARK (Sun, Zhu and Zhou, 2020), BOOST-GP (Li et al., 2021), and GPcounts (BinTayyash et al., 2021) also take advantage of GP regression but directly model raw count data with Poisson, Negative Binomial (NB), and zero-inflated NB (ZINB) distributions, respectively. CTSV (Yu and Luo, 2022), on the other hand, implements a slightly different technique based on ZINB regression, where the mean expression level is parameterized as a linear combination of functions of spatial coordinates. There are some non-parametric approaches with more computational efficiency, such as SPARK-X (Zhu, Sun and Zhou, 2021) based on covariance-based testing, MERINGUE (Miller et al., 2021) based on Voronoi tessellation and classical Moran’s I score, and HEARTSVG (Yuan et al., 2024) based on constant variance testing. Despite considerable successes, the results of the aforementioned works are still sometimes unsatisfactory due to the high dimensionality of genes, high levels of noise and sparsity, and low resolution of spots. In addition to the transcriptomic data, other biological information, such as cellular phenotypes and genetic interactions, is often available and potentially provides a valuable complement to the present analysis. The integration of such assisted information into transcriptomic analysis is a promising direction to mitigate the aforementioned challenges in ST data and further advance existing ST research.
@@
Specifically, first, most existing approaches for SV gene detection rely on marginal analysis, which models each gene separately and has the limitation that the dependence among genes is not well utilized. Increasing evidence has shown that diseases are mostly a result of a combination of multiple genetic alterations, and genetic factors usually interact with each other and are involved in a biological network (Barrio-Hernandez et al., 2023) (Figure 1). Specifically, genes connected within a network are believed to have similar biological functions, leading to potentially similar contributions to cellular organizations and functional mechanisms. In the context of SV gene, many detected SV genes are confirmed to be associated with certain common pathways or networks. For example, the SV genes detected in a human colorectal cancer (CRC) dataset have been found to be enriched in immune-associated GO terms and KEGG pathways (Yuan et al., 2024). Similarly, SV genes identified in human lung and kidney cancer datasets have been confirmed to share certain common biological functionalities (Shang, Wu and Zhou, 2025). Recently, an increasing number of biological networks have been amassed, such as protein-protein interaction (PPI) networks, metabolic networks, and regulatory networks. Curated network information has been widely adopted as a powerful supplement to gene expression analysis, particularly in bulk and single-cell sequencing analysis (Li and Li, 2010; Elyanow et al., 2020; Qin, Ma and Wu, 2023). However, the integration of network information for SV gene detection remains limited.
@@
Second, despite prosperous developments in recent years, measurements obtained using the sequencing-based ST technologies, such as Slide-seq and 10X Genomics Visium, are still “spot”-based, where the gene expression measurement at a single spot is usually a mixture of diverse cells from heterogeneous types rather than at a single-cell resolution (Yu and Luo, 2022) (Figure 1). This cellular composition diversity among spots has been demonstrated to probably contribute to expression variability (Cable et al., 2022a), potentially confounding the detection of biologically relevant SV genes. In the published studies, there are very few approaches that conduct simultaneous cellular composition accommodation and SV gene detection. Limited existing studies include SPARK-X, which converts the inferred major cell types into binary indicators to mitigate the confounding effect caused by latent cell type distributions. Alternatively, CTSV concentrates on the detection of cell-type-specific SV genes whose expressions are affected by the spatial coordinates of cells of the same type.
@@
Motivated by the aforementioned challenges, we propose a Bayesian regularization approach for the joint identification of SV genes as shown in Figure 1. Specifically, the ZINB distribution is adopted to account for the count nature, sparsity, and overdispersion of raw expression measurements. The mean-value-based strategy is applied for SV gene detection, which enjoys the advantages of simplicity, interpretability, and computational efficiency. The most substantial advancement is that the proposed approach conducts joint detection with the network dependency structure among genes well incorporated, making it a big step forward from the existing marginal analysis. Moreover, the proposed approach introduces a series of cell-type-specific parameters to effectively correct for the confounding variations induced by varying cell-type compositions across spots. Overall, this study provides a practically useful and biologically meaningful approach for SV gene identification in ST analysis, with improved performance over alternatives as demonstrated in both simulation studies and application to multiple real-world ST datasets.
Figure 1: Overview of the proposed network-assisted Bayesian regularization framework. Left: ST data consisting of spatial coordinates and corresponding gene expression matrix. Middle: Biological network and cellular composition information for integration. Right: Mean-value-based sparsity-inducing strategy for joint identification of SV genes.
@@ 2.1 Network and Cellular Composition Information for Integration
Consider an undirected network $G(V,E)$ that is constructed using biological information, where $V$ is the node set consisting of $p$ genes and $E=\left\{e(j,l),j,l\in\{1,\right$. $\left.\cdots,p\}\right\}$ is the set of edges between nodes. For genes connected within the network, it is expected that they have similar biological functionalities, leading to potentially similar spatial variability, and thus are more likely to be SV or non-SV genes simultaneously. To induce such network-assisted identification, an adjacency matrix $\boldsymbol{A}=(a_{jl})_{p\times p}$ is first constructed based on $G(V,E)$, with $a_{jl}=1$ if there is an edge $e(j,l)$ between the $j$th and $l$th genes and $a_{jl}=0$ otherwise, and ${a}_{jj}=0$, for $j,l=1,2,\ldots,p$.
As for cellular compositions, denote $\boldsymbol{w}_{i}=\left(w_{i1},\dots,w_{iK}\right)^{\mathrm{T}}$ as the vector of cellular proportions for the $i$th spot, which satisfies the constraint that $0\leq w_{ik}\leq 1,k=1,\dots,K$ and $\sum_{k=1}^{K}w_{ik}=1$. Such information is typically available as ground truth or can be obtained using deconvolution methods such as RCTD (Cable et al., 2022a), Redeconve (Zhou et al., 2023), and SONAR (Liu et al., 2023). Since the distributions of cellular proportions usually display spatial relatedness, which may confound SV gene detection, we correct for this potential confounding in SV gene identification.
@@ 2.2 Network-assisted Bayesian Modeling
Figure 2: Illustrative spatial expression patterns of SV gene with (A) linear, (B) exponential, and (C) periodic pattern with $\beta^{(1)}_{j}\neq 0$ or $\beta^{(2)}_{j}\neq 0$, and (D) non-SV gene with $\beta^{(1)}_{j}=0$ and $\beta^{(2)}_{j}=0$, which are simulated based on the proposed ZINB models (1) and (2).
We adopt the ZINB model due to its superiority for simultaneously accommodating the count measure, over-dispersion, and excessive zeros caused by dropouts. For SV gene identification, significantly different from the previous studies that introduce spatial differential expression via a covariance matrix, we adopt the mean-value-based formulation. Specifically, we introduce the spatialization functions $\mathcal{K}(x_{id})$’s and the corresponding coefficients $\beta^{(d)}_{j}$’s in the mean parameter of the NB distribution to describe the spatial variability. When $\beta^{(1)}_{j}\neq 0$ or $\beta^{(2)}_{j}\neq 0$, the $j$th gene has the expression pattern associated with its spatial location $(x_{i1},x_{i2})$ and is regarded as SV. Such a mean-value based strategy was first introduced by SPARK-X (Zhu, Sun and Zhou, 2021) and has demonstrated strong empirical performance in Yu and Luo (2022); Cable et al. (2022b) and Yu and Li (2024). Compared to covariance-based approaches, our mean-based formulation is more intuitive and interpretable, enabling insights into axis-specific spatial effects about whether expression varies along one specified spatial direction while remaining constant along the other, as shown in Supplementary Figure S13 (Wu et al., 2025). More importantly, the covariance-based strategy always involves constructing and inverting large covariance matrices, which makes the covariance-based strategy usually space- and time- consuming, and thus it lacks efficiency and scalability for large-scale analysis. Here, we follow the published studies (Svensson, Teichmann and Stegle, 2018; Sun, Zhu and Zhou, 2020; Yu and Luo, 2022; Seal, Bitler and Ghosh, 2023) and consider three of the most popular and widely adopted kernel functions to account for a variety of potential spatial patterns that have been discovered in common biological datasets. Illustrative examples are shown in Figure 2 and the detailed settings are provided in Section 2.4. Specifically, a linear function is adopted to recognize a linearly varying expression pattern (Figure 2(A)), while an exponential function is adopted for the expression pattern clustered in focal areas (Figure 2(B)), and a periodic function is particularly useful to detect an expression pattern that is periodically expressed across the spatial area (Figure 2(C)). For the non-SV gene with $\beta^{(1)}_{j}=0$ and $\beta^{(2)}_{j}=0$ (Figure 2(D)), it is observed that the expression pattern is random without association with the spatial location. This mean-value-based strategy requires less storage space and also involves a simpler estimation procedure.
Moreover, we innovatively utilize a set of cell-type-specific $\alpha_{k}$’s for eliminating the confounding impact of the latent cellular composition, where spots that are spatially closer are often observed to have similar cell-type proportions (Cable et al., 2022a). Different from the mean-value-based strategy adopted by Yu and Luo (2022) for marginal cell-type-specific SV gene detection, we focus on the detection of global SV genes while accommodating the cell-type proportion confounding.
@@ 2.3 Priors Specification
where $\boldsymbol{\beta}=\left((\boldsymbol{\beta}^{(1)})^{\mathrm{T}},(\boldsymbol{\beta}^{(2)})^{\mathrm{T}}\right)_{(2p)}^{\mathrm{T}}=\left(\beta^{(1)}_{1},\ldots,\beta^{(1)}_{p},\beta^{(2)}_{1},\ldots,\beta^{(2)}_{p}\right)^{\mathrm{T}}$, $\boldsymbol{\gamma}=\left((\boldsymbol{\gamma}^{(1)})^{\mathrm{T}},(\boldsymbol{\gamma}^{(2)})^{\mathrm{T}}\right)_{(2p)}^{\mathrm{T}}$ represents the effect size of the genes, and $\circ$ denotes the element-wise product. $\mathbf{t}_{\lambda,\boldsymbol{\rho}}(\boldsymbol{\gamma})=\left\{\operatorname{I}\left(\left|\gamma^{(1)}_{1}\right|>\lambda\cdot\rho^{(1)}_{1}\right),\ldots,\operatorname{I}\left(\left|\gamma^{(2)}_{p}\right|>\lambda\cdot\rho^{(2)}_{p}\right)\right\}^{\mathrm{T}}$ is a vector thresholding function with $\lambda$ being a parameter controlling model sparsity and $\boldsymbol{\rho}=\left((\boldsymbol{\rho}^{(1)})^{\mathrm{T}},(\boldsymbol{\rho}^{(2)})^{\mathrm{T}}\right)_{(2p)}^{\mathrm{T}}$ being the adaptive weights. Moreover, $\mathbf{L}=\left(\operatorname{sgn}(\tilde{\boldsymbol{\beta}})\operatorname{sgn}(\tilde{\boldsymbol{\beta}})^{\mathrm{T}}\right)\,\circ\,\widetilde{\mathbf{L}}$ is a block diagonal adaptive Laplacian matrix, where $\operatorname{sgn}(\tilde{\boldsymbol{\beta}})=\left({\operatorname{sgn}}\left(\tilde{\beta}^{(1)}_{1}\right),\ldots,{\operatorname{sgn}}\left(\tilde{\beta}^{(2)}_{p}\right)\right)^{\mathrm{T}}$ with $\tilde{\boldsymbol{\beta}}$ being a rough estimate of $\boldsymbol{\beta}$ and $\widetilde{\mathbf{L}}=\mathbf{I}-\widetilde{\boldsymbol{D}}^{-1/2}\widetilde{\boldsymbol{A}}\widetilde{\boldsymbol{D}}^{-1/2}$ with $\widetilde{\boldsymbol{A}}=\left(\begin{array}[]{cc}\boldsymbol{A}&\mathbf{0}\\ \mathbf{0}&\boldsymbol{A}\end{array}\right)=(\widetilde{a}_{jl})_{2p\times 2p}$ and $\widetilde{\boldsymbol{D}}=\operatorname{diag}\left(d_{1},\cdots,d_{p},d_{1},\cdots,d_{p}\right)$ with $d_{j}=\sum_{l=1}^{p}a_{jl}$. $\varepsilon$ is a small constant to make ${\mathbf{L}}+\varepsilon\mathbf{I}_{2p}$ strictly positive-definite, which is set as $10^{-3}$ in our numerical work.
@@
The proposed priors have been motivated by the following considerations. The identification of SV genes is achieved using the hard-thresholded Gaussian prior, where $\beta^{(d)}_{j}$ is shrunk to zero when $\mathrm{t}_{\lambda,\rho^{(d)}_{j}}\left(\gamma^{(d)}_{j}\right)=\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\cdot\rho^{(d)}_{j}\right)=0$ ($d=1,2$). The thresholded Gaussian prior has been shown as a useful alternative to shrinkage priors in Bayesian sparse analysis (Wu, Guo and Kang, 2024), and is favored here for its simplicity and flexibility as well as its appealing interpretability as a minimal detectable signal. Here, we further introduce a series of weights $\rho_{j}^{(d)}$’s in $\mathbf{t}_{\lambda,\boldsymbol{\rho}}(\boldsymbol{\gamma})$ to adjust the shrinkage of various $\gamma^{(d)}_{j}$’s to improve selection efficiency, where the genes with strong spatial variability are potentially assigned small weights and thus are more likely to have nonzero $\beta^{(d)}_{j}$’s.
@@
Moreover, the network dependency is introduced via the graph Laplacian matrix in the covariance matrix of the hard-thresholded Gaussian prior. The proposed network-assisted strategy is motivated by the successes of Laplacian shrinkage in high-dimensional regression analysis (Chakraborty and Lozano, 2019; Cai, Kang and Yu, 2020). Different from these studies, we innovatively conduct SV gene detection with the network structures among genes incorporated. In particular, the Laplacian matrix $\widetilde{\mathbf{L}}$ is further modified with a pre-defined sign matrix $\operatorname{sgn}(\tilde{\boldsymbol{\beta}})\operatorname{sgn}(\tilde{\boldsymbol{\beta}})^{\mathrm{T}}$ to accommodate the scenario where two neighborhood genes are negatively correlated and have opposite directions of spatial variability. With the proposed priors, we have $\boldsymbol{\gamma}^{\mathrm{T}}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)\boldsymbol{\gamma}=\sum_{j\sim l}\left(\frac{\text{sgn}\left(\tilde{\beta}_{j}^{(1)}\right)\gamma_{j}^{(1)}}{\sqrt{d_{j}}}-\frac{\text{sgn}\left(\tilde{\beta}_{l}^{(1)}\right)\gamma_{l}^{(1)}}{\sqrt{d_{l}}}\right)^{2}+\sum_{j\sim l}\left(\frac{\text{sgn}\left(\tilde{\beta}_{j}^{(2)}\right)\gamma_{j}^{(2)}}{\sqrt{d_{j}}}-\frac{\text{sgn}\left(\tilde{\beta}_{l}^{(2)}\right)\gamma_{l}^{(2)}}{\sqrt{d_{l}}}\right)^{2}$, where for genes $j$ and $l$ with an edge $e(j,l)$, the absolute values of $\gamma^{(d)}_{j}$ and $\gamma^{(d)}_{l}$ are promoted to be similar, further inducing simultaneous SV or non-SV.